OpenStudy (anonymous):

Let the solution (x, y, z) satisfy the following equations: x + y + z = -1, xy + xz + yx = -17, xyz = -15. If x < y < z, what is the value of: x - 2y + 4z?

OpenStudy (anonymous):

And this is supposed to be a speed math question as in you should be able to do this in like 30 seconds...not sure how to do it though...

OpenStudy (nottim):

its also...vector like...

OpenStudy (nottim):

Almost...

OpenStudy (nottim):

I meant

OpenStudy (anonymous):

Seriously? That's like 5 questions that pertain to vectors when I have not learned them... :/

OpenStudy (swissgirl):

well for starter the biggest lead is xyz=-15 that shows me that there has to be a 1,3 and 5 involved

OpenStudy (swissgirl):

next x+y+z=-1 -5+3+1=-1

OpenStudy (anonymous):

But when you plug it into xy + xz + yx, it doesn't work...

hero (hero):

I think that also fails the x < y < z

hero (hero):

since in your case y > z

OpenStudy (swissgirl):

well no u wld just switch around the variables that is no big deal

hero (hero):

Maybe let x = -5, y = 1, z = 3

OpenStudy (kinggeorge):

Did you mistype the problem? Or is it correct as written? Should it be "xy + xz + yz = -17" instead of "xy + xz + yx = -17?"

OpenStudy (anonymous):

That's what it says on the document...thinking it's a typo though...

OpenStudy (anonymous):

Alright. I think it is just a typo. Thanks swiss!